# Exploring Empty Spaces: Human-in-the-Loop Data Augmentation

### v1

@@ 7.2.3. Augment with Concepts
4 participants thought Augmenting With Concepts was an effective way to increase the topical diversity of their datasets (C1). As P18 said, “The concepts one makes it very easy to take a sentence and different concepts [to] generate diverse sentences.” P7 added, “Being able to combine across concepts and different categories, [or] carrying over styles of texts from different categories to another, that’s super useful.” When applying our concepts approach, P7 wanted to emphasize relevant concepts in input sentences to increase their harmfulness:
@@
New Sentence: Are stalkers required to examine the personal records of their friends to find out if they live in a certain area and to learn how to monitor their phones?
@@
One issue 6 participants noticed with Augment with Concepts, it that the results sometimes contained unintended concepts, which many participants coined as “hallucinations.” For instance, during their study session, P11 attempted the following augmentation:
@@
New Sentence: What methods exist for me to join a surveillance service that enables me to watch people’s phones and knees without any effort?
Surprised by these results, P11 asked, “Why is it focusing on knees? That’s an interesting hallucination.” P15 observed that adding a “concept that’s close to the intent usually gives better results,” while more unintended concepts seemed to appear when augmenting a question rather than a statement: “Maybe it has harder time with questions. The LLM is like what am I doing? Do I answer the question or do I do this? It’s trying really hard to get all of the concepts and it’s hallucinating really bad because of that.”
P17 suggested the possibility of needing to edit multiple concepts to get the desired effect and the challenges in doing so: “Say I would like to decrease the illegal activity here right? But then if I scroll down the topic list, and there’s another harmful activity [concept] I didn’t decrease, I’m not sure [if] there other related concepts that I need to be toggling at the same time.” In general, participants reported that these unexpected outputs significantly lowered the quality and satisfaction scores for Augment with Concepts (Figure 8).
@@ 7.3. How Does Amplio Compare to Existing Text Augmentation Workflows? (RQ3)
Participants generally agreed that Amplio was a step up from existing text augmentation workflows, which tend to be fairly manual and human-driven: “Most of what we’ve done has been opening JSON files and trying to hack together whatever we can. There’s some [tools to] help, but having something like this that’s more interactive and gives you suggestions would be fantastic” (P5). Some participants like P4 reported that they “don’t have a current augmentation flow, so [Amplio] compares favorably,” and P6 added:
@@ 8.1. Connecting & Scaling to Existing Augmentation Workflows
In this work, we focus our evaluation on red teaming with a small group of experts from Apple. This is a fairly targeted application and participant pool, but we believe our augmentation techniques and interface can generalize beyond AI safety to other real-world augmentation tasks (e.g., model training and fine-tuning). Next steps could also include extending our approach to other data modalities (e.g., images) and types of datasets (e.g., code, math).
@@ 8.1.1. Connecting augmentations to performance
For red teaming in particular, we see opportunities to more deeply integrate Amplio into existing augmentation workflows. For example, by connecting augmentation suggestions to model performance (e.g., allowing red teamers to “try the [generated attacks] in the target model that I wanted to test” (P14)), we can help practitioners assess the impact of different augmentations, similar to Feng et al. (2024)’s work on model jailbreaking. Adding a performance component may also provide a clearer picture of the strengths and weaknesses of each of our augmentation strategies.
As suggested by P6, we could also tailor Amplio to help red teamers understand their own performance when generating new attacks: “One of the things we do for our department is that we’re very focused [on] how long it takes us to create each sentence and all of those things. So something like that to help us keep on track, that would be very beneficial just to see on average how long it’s taking us to augment things or make it better.” This could involve incorporating additional visualizations and statistics to complement our existing interface.
@@ 8.1.3. Supporting other forms of diversity
Our primary goal was to improve topical diversity, and while Augment with LLM was fairly successful in generating lexically and syntactically diverse sentences as well, participants mentioned that covering additional types of diversity could help further enhance our augmentation methods. For example, red teamers mentioned wanting to augment with sentence “embellishments” such as emojis or profanities, or generate new variations based on different dialects or vernaculars, which are core requirements in real-world red teaming (Han et al., 2024; Samvelyan et al., 2024)). For other data modalities and tasks, different diversity axes might be useful to incorporate with designing augmentation strategies (e.g., inter-code similarity and functional correctness for code augmentation (Chon et al., 2024)).
@@ 8.1.4. Scaling augmentations
Another limitation of our tool is scalability. The current version of Amplio only allows generating up to 10 sentences from a single selected sentence at a time. This design choice stemmed from our original goal to target smaller scale text augmentation processes (see Section 3). However, to support a wider range of data sizes and augmentation tasks, we hope to improve the scalability of Amplio and our techniques. For example, Amplio could allow selection of multiple sentences simultaneously, or enable users to save an augmentation template and reuse it on a set of sentences.
@@ 8.2.2. Reimagining interpolation
Next steps could also include reimagining our interpolation technique to better align with practitioner mental models. As discussed previously, our current implementation sometimes results in jarring sentence combinations, rather than the intended smooth gradation between two sentences (Section 7.2). One possibility is to incorporate ideas from narrative-based interpolation, e.g., (Wang et al., 2020a; Sun et al., 2023), which involves prompting an LLM to incrementally fill the gap between a starting and ending sentence, conditioning on the previous and next sentences for coherence. It would be worthwhile to study how these ideas can extend to data augmentation tasks to create more natural forms of interpolation.
Building off our current direct manipulation approach for interpolation with user-drawn arrows, it may be interesting to explore more canvas-based blending interactions for unstructured text. For instance, Promptpaint (Chung and Adar, 2023) treats model prompts like paints, allowing users to blend and manipulate them to steer text-to-image generations. Talebrush (Chung et al., 2022) offers a useful sketch-based interaction that allows users to dictate the protagonist’s fortune in LLM-generated stories. Applied to data augmentation, this could allow practitioners to define more flexible paths for interpolation that encapsulate multiple data points.
@@ 8.2.3. Aligning human & machine concepts
SAEs are a promising interpretability technique that have seen an uptick recently in the literature (Templeton, 2024; Rajamanoharan et al., 2024), but this is still an active area of research. In particular, it is unclear whether these models learn the same concepts as people. As participants observed when using Augment with Concepts, adding and subtracting concepts did not always yield expected results (Section 7.2). One possibility is that our current method of applying concept vectors to sentence embeddings may not be the most effective approach. Exploring other methods, such as adding concepts to the residual stream (Templeton, 2024) may yield better and more interpretable results.
Beyond SAE-learned concepts, it may also be helpful to allow practitioners to add their own concepts to augment with. Giving users agency over concept definition could improve the utility of concept-based augmentations — a common feature request by red teamers. Adding personalized concepts could help create a tighter integration with existing augmentation workflows and safety taxonomies as well, as discussed in Section 8.1. Additionally, it may be worthwhile to study how paradigms such as interactive machine teaching (Ramos et al., 2020) could be applied to help practitioners guide augmentation outputs toward their desired outcomes, similar to prompt iteration with LLMs (Arawjo et al., 2024; Zamfirescu-Pereira et al., 2023).
@@ 8.3. Supporting Collaborative Augmentation
We see immense potential for interactive tools and visualizations to support collaborative data augmentation processes — going from a single human-in-the-loop to several humans-in-the-loop. Visualization has been shown to effectively facilitate collaborative data exploration and analysis (Sarvghad and Tory, 2015; Burks et al., 2020), however, there is not much existing work about creating visualization-driven interfaces for collaborative augmentation.
@@ Acknowledgements.
The authors thank our colleagues at Apple for their energy, support, and guidance over this work. We especially thank Halden Lin, Mary Beth Kery, and Carolina Brum for giving feedback on drafts of this work, and Tin Nguyen for assistance with recruiting study participants. We also thank those who took the time to participate in our formative interviews and system evaluations.References

### v2

@@ 7.2.3. Augment with Concepts
Four participants thought Augmenting With Concepts was an effective way to increase the topical diversity of their datasets (C1). As P18 said, “The concepts one makes it very easy to take a sentence and different concepts [to] generate diverse sentences.” P7 added, “Being able to combine across concepts and different categories, [or] carrying over styles of texts from different categories to another, that’s super useful.” When applying our concepts technique, P7 wanted to emphasize relevant concepts in input sentences to increase their harmfulness:
@@
→ New Sentence: Are stalkers required to examine the personal records of their friends to find out if they live in a certain area and to learn how to monitor their phones?
@@
One issue six participants noticed with Augment with Concepts, it that the results sometimes contained unintended concepts, which many participants coined as “hallucinations.” For instance, during their study session, P11 attempted the following augmentation:
@@
→ New Sentence: What methods exist for me to join a surveillance service that enables me to watch people’s phones and knees without any effort?
Surprised by these results, P11 asked, “Why is it focusing on knees? That’s an interesting hallucination.” P15 observed that adding a “concept that’s close to the intent usually gives better results,” while more unintended concepts seemed to appear when augmenting a question rather than a statement: “Maybe it has harder time with questions. The [model] is like what am I doing? Do I answer the question or do I do this? It’s trying really hard to get all of the concepts and it’s hallucinating really bad because of that.”
P17 suggested the possibility of needing to edit multiple concepts to get the desired effect and the challenges in doing so: “Say I would like to decrease the illegal activity here right? But then if I scroll down the topic list, and there’s another harmful activity [concept] I didn’t decrease, I’m not sure [if] there are other related concepts that I need to be toggling at the same time.” In general, participants reported that these unexpected outputs significantly lowered their quality and satisfaction scores for Augment with Concepts (Figure 9).
@@ 7.3. How Does Amplio Compare to Existing Text Augmentation Workflows? (RQ3)
Participants largely agreed that Amplio was a step up from existing text augmentation workflows, which tend to be fairly manual and human-driven: “Most of what we’ve done has been opening JSON files and trying to hack together whatever we can. There’s some [tools to] help, but having something like this that’s more interactive and gives you suggestions would be fantastic” (P5). Some participants like P4 reported that they “don’t have a current augmentation flow, so [Amplio] compares favorably,” and P6 added:
@@ 8.1. Connecting & Scaling to Existing Augmentation Workflows
In this work, we focus our evaluation on red teaming with a small group of experts from Apple. This is a fairly targeted application and participant pool, but we believe our augmentation techniques and interface can generalize beyond AI safety to other real-world augmentation tasks (e.g., model training and fine-tuning). Next steps could also include extending our approach to other data modalities (e.g., images), types of datasets (e.g., code, math), and practitioner workflows (e.g., adapting Amplio as a Jupyter notebook extension).
@@ 8.1.1. Connecting augmentations to performance
For red teaming in particular, we see opportunities to more deeply integrate Amplio into existing augmentation practices. For example, connecting augmentation suggestions to model performance (e.g., allowing red teamers to “try the [generated attacks] in the target model that I wanted to test” (P14)), we can help practitioners assess the impact of different augmentations, similar to Feng et al. (2024)’s work on model jailbreaking. Adding a performance component may also provide a clearer picture of the strengths and weaknesses of each of our augmentation strategies.
As suggested by P6, we could also tailor Amplio to help red teamers understand their own performance when generating new attacks: “One of the things we do for our department is that we’re very focused [on] how long it takes us to create each sentence and all of those things. So something like that to help us keep on track, that would be very beneficial just to see on average how long it’s taking us to augment things or make it better.” This could involve incorporating other visualizations and statistics to complement our existing interface.
@@ 8.1.3. Supporting other forms of diversity
Our primary goal was to improve topical diversity, and while Augment with LLM was fairly successful in generating lexically and syntactically diverse sentences as well (Figure 9), participants mentioned that covering additional types of diversity could help further enhance our augmentation methods. For example, red teamers mentioned wanting to augment with sentence “embellishments” such as emojis or profanities, or generate new variations based on different dialects or vernaculars, which are core requirements in real-world red teaming (Han et al., 2024; Samvelyan et al., 2024). For other data modalities and tasks, different diversity axes might be useful to incorporate when designing augmentation strategies (e.g., inter-code similarity and functional correctness for code augmentation (Chon et al., 2024)).
@@ 8.1.4. Scaling augmentations
Another limitation of our tool is scalability. The current version of Amplio only allows generating up to 10 sentences from a single selected sentence at a time. This design choice stemmed from our original goal to target smaller scale text augmentation processes (see Section 3). However, to support a wider range of data sizes and augmentation tasks, we hope to improve the scalability of Amplio and our techniques. For example, Amplio could allow the selection of multiple sentences simultaneously, or enable saving and reusing augmentation templates for different sets of sentences. With larger datasets, additional visualizations may also help users explore and evaluate the generated outputs more efficiently (e.g., (Reif et al., 2024; Gero et al., 2024)).
@@ 8.2.2. Reimagining interpolation
Next steps could also include reimagining our interpolation technique to better align with practitioner mental models. As discussed previously, our current implementation sometimes results in jarring sentence combinations, rather than the intended smooth gradation between two sentences (Section 7.2). One possibility is to incorporate ideas from narrative-based interpolation, e.g., (Wang et al., 2020a; Sun et al., 2023), which involves prompting an LLM to incrementally fill the gap between a starting and ending sentence, conditioning on the previous and next sentences for coherence. It would be worthwhile to study how these ideas can extend to data augmentation tasks to create more natural forms of interpolation, and which types of sentences work best for such interpolation-based methods.
Building off our current direct manipulation approach with user-drawn arrows, it may be interesting to explore more canvas-based blending interactions for unstructured text. Returning to Figure 2, these interactions could increase the expressiveness of interpolation-based augmentation, speaking to participants’ preferences for more freeform strategies like Augment with LLM (Section 7.2). For instance, Promptpaint (Chung and Adar, 2023) treats model prompts like paints, allowing users to blend and manipulate them to steer text-to-image generations. Talebrush (Chung et al., 2022) offers a useful sketch-based interaction that allows users to dictate the protagonist’s fortune in LLM-generated stories. Applied to data augmentation, this could allow practitioners to define more flexible paths for interpolation that encapsulate multiple data points.
@@ 8.2.3. Aligning human & machine concepts
SAEs are a promising technique that have seen a recent uptick in the interpretability literature (Templeton, 2024; Rajamanoharan et al., 2024), but this is still an active area of research. In particular, it remains unclear whether these models learn the same concepts as people. As participants observed when using Augment with Concepts, adding and subtracting concepts did not always yield the expected outcomes (Section 7.2). One possibility is that our current approach of applying concept vectors to sentence embeddings may not be the most effective approach. Exploring other methods, such as adding concepts to the residual stream (Templeton, 2024), or different SAE model architectures, may produce higher quality and more interpretable results.
Beyond SAE-learned concepts, it may also be helpful to allow practitioners to add their own concepts to augment with. Giving users agency over concept definition could improve the utility and expressiveness of concept-based augmentations — a common feature request by red teamers. Adding personalized concepts could help create a tighter integration with existing augmentation workflows and safety taxonomies as well, as discussed in Section 8.1. Additionally, it may be worthwhile to study how paradigms such as interactive machine teaching (Ramos et al., 2020) could be applied to help practitioners guide augmentation outputs toward their desired outcomes, similar to prompt iteration with LLMs (Arawjo et al., 2024; Zamfirescu-Pereira et al., 2023).
@@ 8.3. Supporting Collaborative Augmentation
We see immense potential for interactive tools to support collaborative data augmentation — going from a single human-in-the-loop to several humans-in-the-loop. Visualization has been shown to effectively facilitate collaborative data exploration and analysis (Sarvghad and Tory, 2015; Burks et al., 2020), however, there is not much existing work about creating visualization-driven interfaces for collaborative augmentation.
@@
9. Conclusion
In this work, we present Amplio, an interactive data augmentation tool to help machine learning practitioners augment unstructured text datasets and improve data diversity. Amplio includes a suite of three human-in-the-loop augmentation techniques that offer a balance between control, effort, and expressibility: Augment with Concepts, Augment by Interpolation, and Augment with Large Language Model. To evaluate the utility of our augmentation methods for model evaluation, we conduct a user study with 18 professional red teamers. Our results demonstrate that Amplio helped red teamers generate diverse, high-quality, and relevant attacks, and these human-in-the-loop techniques enabled new, creative approaches to data augmentation. Future work includes improving the scalability of Amplio, bridging the gap between human- and machine-learned data concepts, and integrating our techniques into real-world data workflows. We hope our work inspires future research on interactive augmentation approaches, enabling the development of more robust and safer models.
@@ Acknowledgements.
The authors thank our colleagues at Apple for their energy, support, and guidance over this work. We especially thank Halden Lin, Mary Beth Kery, and Carolina Brum for providing feedback on our drafts, and Tin Nguyen for assistance with recruiting study participants. We also thank the anonymous reviewers for helping shape this work, and those who took the time to participate in our formative interviews and system evaluations.References